OpenStudy (anonymous):

[PLEASE HELP - I've been waiting 40 minutes. I will give you a medal if you help. Please help me.] Suppose the maximum healthy weight for a person who is 5 feet 9 inches tall is 150 pounds, and the maximum healthy weight for someone 6 feet 7 inches tall is 200 pounds. The relationship between weight and height here is linear. (a) Find a linear equation that gives the maximum healthy weight y for a person whose height is x inches over 4 feet 9 inches. (Thus x = 0 corresponds to 4 feet 9 inches and x = 3 to 5 feet,etc.)

OpenStudy (anonymous):

i cannot work with feet, inches so lets convert to inches

OpenStudy (anonymous):

OK, x=0 equals 4' 9" and x=3 equals 5'

OpenStudy (anonymous):

five foot nine is 69 inches six foot seven is 79 inches

OpenStudy (anonymous):

four foot nine is 57 inches

OpenStudy (anonymous):

so we subtract

OpenStudy (anonymous):

No, but those values need to be relative to 4' 9". I need the equation that gives the maximum healthy weight y for a person whose height is x inches OVER 4 feet 9 inches.

OpenStudy (anonymous):

\[69-57=12,79-57=22\]

OpenStudy (anonymous):

find the equation of the line through the points \((12,150)\) and \((22,200)\) that is all

OpenStudy (anonymous):

you cannot work in math with units like so many feet so many inches you have to have one unit of measurement to find an equation like this

OpenStudy (anonymous):

you can almost do it in your head the slope is \(\frac{50}{10}=5\)